1. There are 49 dogs signed up for a dog show. There are 36 more small dogs than large dogs. How many small dogs have signed up to compete? 

This question comes directly from a second grader's math homework.

1
Expert's answer
2022-06-02T17:24:45-0400

If there are x large dogs then small dogs are x + 36 and together there are 49 dogs. This sentence can be rewritten in the equation:

x + x + 36 = 49.

Let's find the solution.

2x + 36 = 49;

2x = 13;

x = 6.5 - number of large dogs,

and x + 36 = 6.5 + 36 = 42.5 - number of small dogs.

Nevertheless, the number of dogs should be integer. It is impossible there are medium dogs. The question seems to have incorrect conditions.